∠A and ∠B are supplementary. If m∠A = 6x - 1
and m∠B = 5x - 17, find both measures.

Answer by ikleyn(52869)

.
∠A and ∠B are supplementary. If m∠A = 6x - 1
and m∠B = 5x - 17, find both measures.
~~~~~~~~~~~~~~~~~~~~~~~~~~

The sum of supplementary angles is 180 degrees

    (6x-1) + (5x-17) = 180  degrees

    6x - 1 + 5x - 17 = 180

    6x + 5x = 180 + 1 + 17

      11x   =     198

        x   =     198/11 = 18


m∠A = 6x-1 = 6*18-1 = 107 degrees;  m∠B = 5*18-17 = 73 degrees.    ANSWER

Solved.